OpenStudy (anonymous):

Which of the following is true about the function below? (will draw in a sec!) (and will also post answer choices in a sec!!) ***PLease explain!! :) thanks :D

OpenStudy (anonymous):

 this is the function :)

OpenStudy (anonymous):

asnwers A,B,C,D (top to bottom)

OpenStudy (anonymous):

Hi!! Thanks for coming!! :) DO u get this?? I am uber confused :(

jimthompson5910 (jim_thompson5910):

you can't divide by zero, so \(\Large \sqrt{x-6} \neq 0\)

OpenStudy (anonymous):

Well, it's domain is when \( \sqrt{x-6} \neq 0 \) (can't divide by 0) and \(x-6 \geq 0\) (no negative square roots). It's range must be above 0, since it can't ever be negative.

jimthompson5910 (jim_thompson5910):

If the denominator were zero, then \(\Large \sqrt{x-6} = 0\) \(\Large x-6 = 0^2\) \(\Large x-6 = 0\) \(\Large x = 0+6\) \(\Large x = 6\) so \(\Large x = 6\) makes the denominator zero, so \(\Large x \neq 6\) to avoid division by zero

OpenStudy (anonymous):

so how do i find the range and domain tho? i don't get that :(

OpenStudy (anonymous):

so x can't equal 6 and it can't equal sq. rt. of x-6 ?? did i get that right?

jimthompson5910 (jim_thompson5910):

also, you can't take the square root of a negative number, so the radicand is positive (it's not zero for reasons stated above) \(\Large x - 6 > 0\) \(\Large x > 0+6\) \(\Large x > 6\) So the domain is \(\Large x > 6\) which in interval notation is \(\Large \left(6,\infty\right)\)

OpenStudy (anonymous):

oh i get how to do that now! :) but how do i find the range?

jimthompson5910 (jim_thompson5910):

to find the range, you start with \[\Large y = \frac{1}{\sqrt{x-6}}\] You then swap x and y \[\Large x = \frac{1}{\sqrt{y-6}}\] Then solve for y \[\Large x\sqrt{y-6} = 1\] \[\Large \sqrt{y-6} = \frac{1}{x}\] \[\Large y-6 = \left(\frac{1}{x}\right)^2\] \[\Large y-6 = \frac{1}{x^2}\] \[\Large y = \frac{1}{x^2}+6\] \[\Large y = \frac{1}{x^2}+6*\frac{x^2}{x^2}\] \[\Large y = \frac{1}{x^2}+\frac{6x^2}{x^2}\] \[\Large y = \frac{1+6x^2}{x^2}\] Now find the domain of the last equation. This will give you the range of the original function.

OpenStudy (anonymous):

y=7?

jimthompson5910 (jim_thompson5910):

you can't divide by zero right?

OpenStudy (anonymous):

no :/

OpenStudy (anonymous):

but the only answer choices left that would work are A and C right?  with parentheses right? not brackets??